# SAGE: An Agentic Explainer Framework for Interpreting SAE Features in Language Models

### v1

@@ Abstract
Jiaojiao Han1   Wujiang Xu2   Mingyu Jin2   Mengnan Du1 1New Jersey Institute of Technology 2Rutgers University
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress across diverse domains, including natural language understanding, generation, and reasoning. However, despite their impressive capabilities, LLMs remain largely opaque systems, often regarded as black boxes whose internal mechanisms are poorly understood zhao2024explainability. To address this opacity, the research community has increasingly focused on decoding the information encoded in LLM representations, seeking to understand how these models process and store knowledge. Among various interpretability approaches, sparse autoencoders (SAEs) have attracted growing attention due to their ability to decompose dense neural activations into sparse, potentially interpretable features shu2025survey. Recent work has demonstrated that SAEs can identify meaningful feature dimensions in transformer representations, with applications ranging from circuit discovery to activation steering ferrando2024know; he2025sae.
Despite this progress, interpreting SAE features remains a significant challenge. As SAEs are trained using unsupervised learning objectives, the semantic meaning of their learned features must be inferred post-hoc through analysis of their activation patterns. Current approaches, exemplified by Neuronpedia neuronpedia, rely on automated interpretation pipelines that generate natural language explanations for each SAE feature using large language models such as GPT-4 and Claude 4.5. While these methods have produced preliminary results, two fundamental problems persist. First, the generated explanations lack consistency and rigor. When different LLMs are used to explain the same feature, they often produce divergent explanations, undermining confidence in the interpretations. Second, although SAEs are explicitly designed to decompose polysemous LLM representations into monosemantic features, where each feature captures a single, coherent concept. In practice, many SAE features still exhibit polysemantic behavior, activating in response to multiple distinct semantic or structural patterns. Existing methods like Neuronpedia provide only a single explanation per feature, failing to account for this multi-faceted activation behavior and potentially missing important aspects of feature functionality.
@@ 2.1 Sparse Autoencoders
SAEs bricken2023towards; cunningham2023sparse; templeton2024scaling are designed to address the opacity of large models by decomposing dense neural activations $x\in\mathbb{R}^{d_{\text{model}}}$ into sparse, potentially interpretable features $f\in\mathbb{R}^{d_{\text{sae}}}$. This is achieved by projecting the input into a much higher-dimensional feature space, where $d_{\text{sae}}\gg d_{\text{model}}$.The architecture consists of an encoder that computes the sparse features $f$, and a decoder that uses these sparse features to reconstruct the original activation, $\hat{x}$:
@@ Appendix A More Details of the Evaluation Metrics
  • Predictive Accuracy. This metric assesses the descriptive power of an explanation: can it be used to predict feature activations on held-out data? We use a held-out set of exemplars $D_{\text{held-out}}$, distinct from the $D_{j,k}$ used for explanation generation, sampled from high, medium, and low activation groups. Following past work cunningham2023sparse, we employ a simulator $\sigma$, which is an LLM prompted with the feature explanation $E_{j}$. For each token $t$ in a held-out example, $\sigma$ predicts the discretized activation level. Rather than single-point prediction, we compute the expected activation value using the log-probabilities $\sigma$ assigns to the output tokens ’0’ through ’10’. The predictive accuracy is the mean Pearson correlation coefficient ($\rho$) between the predicted activation values and the true, normalized per-token activations across $D_{\text{held-out}}$.
@@ Appendix B Related Work
Sparse autoencoders (SAEs). SAEs were introduced as an unsupervised dictionary-learning approach to address superposition faruqui2015sparse in LLM shu2025survey; huben2023sparse. By mapping model activations into a higher-dimensional sparse space, SAEs isolate a small number of latent features per input, yielding monosemantic features that correspond to single interpretable concepts rather than polysemantic neurons bricken2023monosemanticity. A number of SAE variants and tools have been developed to improve their efficacy and accessibility. The vanilla SAE typically uses an $L_{1}$ sparsity penalty on the latent vector to encourage most neurons to stay inactive sharkey2022sae and recent variants like the Top-$K$ SAE instead enforce a fixed number $K$ of active features per input gao2025scaling. Other improvements include gated or JumpReLU SAEs that modify the activation function to better balance feature detection and strength estimation rajamanoharan2024jumping; rajamanoharan2024improving. Some pre-trained repositories, such as Gemma Scope lieberum2024gemma and Llama Scope he2024llama, enable broader research.
@@
SAEs Application. SAEs have been used to interpret model representations and understand model capabilities wu2025self; wu2025interpreting. Beyond static analysis, researchers have begun leveraging SAE-discovered features to steer model behavior. Such activation steering via SAE features has been used to alter attributes like sentiment, truthfulness, or style without fine-tuning the entire model he2025sae; shi-etal-2025-route; wang-etal-2025-beyond-prompt; wang2025two. SAEs have also been applied in the context of model safety and alignment. One study showed that features learned by an SAE from a language model can serve as effective probes for classifying toxic content across languages bereskamechanistic. By identifying which sparse features correspond to a model’s refusals or safety responses, one can understand and even adjust the model’s safety mechanisms. Intervening on these features has been shown to influence the model’s tendency to refuse or comply with certain prompts arditi2024refusal; yeo-etal-2025-understanding; wu2025automating. Overall, SAEs offer a transparent, feature-level handle on model behaviors that is valuable for safety research.
@@
SAEs Feature Explanation. Inspired by the automated interpretability pipeline that uses GPT-4 to explain GPT-2 neurons from their activating examples (MaxAct) bills2023language, a framework that has since become the standard for large-scale interpretation of neurons and SAE-learned features in both language and vision models neuronpedia; huben2023sparse; gao2025scaling. Neuronpedia combines an activation-based method kissane2024interpreting that highlights the tokens most strongly triggering a feature with a logit-projection method kissane2024interpreting that infers the feature’s semantic direction by measuring its positive and negative influence on output logits. Recent work proposes an "output-centric" automated feature interpretation that interprets model features not only by considering which inputs activate them, but also by examining the impact of their activation on the model output to generate more accurate and causal interpretations gur-arieh-etal-2025-enhancing.
@@
Agents for Explainability. Recent work has explored using agentic frameworks for explainability. For instance, MAIA shaham2024multimodal employs a vision-language model equipped with a set of tools to automate the interpretation of computer vision models. MAIA iteratively designs experiments, composes tools for tasks like input synthesis and exemplar generation, and formulates explanations to explain model behaviors, such as identifying feature selectivity or failure modes. Similarly, KnowThyself prasai2025knowthyself provides an agentic assistant specifically for LLM interpretability. It unifies various interpretability tools into a single chat-based interface, allowing users to ask natural language questions. In contrast to these applications, our work proposes an agent framework specifically designed to interpret the features learned by SAEs.

### v2

@@ Abstract
Jiaojiao Han1   Wujiang Xu2   Mingyu Jin2   Mengnan Du3 1New Jersey Institute of Technology  2Rutgers University 3The Chinese University of Hong Kong, Shenzhen liuliujiujiu05@gmail.commengnandu@cuhk.edu.cn Corresponding author
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress across diverse domains, including natural language understanding, generation, and reasoning. However, despite their impressive capabilities, LLMs remain largely opaque systems, often regarded as black boxes whose internal mechanisms are poorly understood Zhao et al. (2024). To address this opacity, the research community has increasingly focused on decoding the information encoded in LLM representations, seeking to understand how these models process and store knowledge. Among various interpretability approaches, sparse autoencoders (SAEs) have attracted growing attention due to their ability to decompose dense neural activations into sparse, potentially interpretable features Shu et al. (2025). Recent work has demonstrated that SAEs can identify meaningful feature dimensions in transformer representations, with applications ranging from circuit discovery to activation steering Ferrando et al. (2025); He et al. (2025).
Despite this progress, interpreting SAE features remains a significant challenge. As SAEs are trained using unsupervised learning objectives, the semantic meaning of their learned features must be inferred post-hoc through analysis of their activation patterns. Current approaches, exemplified by Neuronpedia Lin (2023), rely on automated interpretation pipelines that generate natural language explanations for each SAE feature using large language models such as GPT-4 and Claude 4.5. While these methods have produced preliminary results, two fundamental problems persist. First, the generated explanations lack consistency and rigor. When different LLMs are used to explain the same feature, they often produce divergent explanations, undermining confidence in the interpretations. Second, although SAEs are explicitly designed to decompose polysemous LLM representations into monosemantic features, where each feature captures a single, coherent concept. In practice, many SAE features still exhibit polysemantic behavior, activating in response to multiple distinct semantic or structural patterns. Existing methods like Neuronpedia provide only a single explanation per feature, failing to account for this multi-faceted activation behavior and potentially missing important aspects of feature functionality.
@@ 2.1 Sparse Autoencoders
SAEs Bricken et al. (2023b); Cunningham et al. (2023); Templeton et al. (2024) are designed to address the opacity of large models by decomposing dense neural activations $x\in\mathbb{R}^{d_{\text{model}}}$ into sparse, potentially interpretable features $f\in\mathbb{R}^{d_{\text{sae}}}$. This is achieved by projecting the input into a much higher-dimensional feature space, where $d_{\text{sae}}\gg d_{\text{model}}$.The architecture consists of an encoder that computes the sparse features $f$, and a decoder that uses these sparse features to reconstruct the original activation, $\hat{x}$:
@@ Appendix A More Details of the Evaluation Metrics
  • Predictive Accuracy. This metric assesses the descriptive power of an explanation: can it be used to predict feature activations on held-out data? We use a held-out set of exemplars $D_{\text{held-out}}$, distinct from the $D_{j,k}$ used for explanation generation, sampled from high, medium, and low activation groups. Following past work Cunningham et al. (2023), we employ a simulator $\sigma$, which is an LLM prompted with the feature explanation $E_{j}$. For each token $t$ in a held-out example, $\sigma$ predicts the discretized activation level. Rather than single-point prediction, we compute the expected activation value using the log-probabilities $\sigma$ assigns to the output tokens ’0’ through ’10’. The predictive accuracy is the mean Pearson correlation coefficient ($\rho$) between the predicted activation values and the true, normalized per-token activations across $D_{\text{held-out}}$.
@@ Appendix B Related Work
Sparse autoencoders (SAEs). SAEs were introduced as an unsupervised dictionary-learning approach to address superposition Faruqui et al. (2015) in LLM Shu et al. (2025); Huben et al. (2024). By mapping model activations into a higher-dimensional sparse space, SAEs isolate a small number of latent features per input, yielding monosemantic features that correspond to single interpretable concepts rather than polysemantic neurons Bricken et al. (2023a). A number of SAE variants and tools have been developed to improve their efficacy and accessibility. The vanilla SAE typically uses an $L_{1}$ sparsity penalty on the latent vector to encourage most neurons to stay inactive Sharkey et al. (2022) and recent variants like the Top-$K$ SAE instead enforce a fixed number $K$ of active features per input Gao et al. (2025). Other improvements include gated or JumpReLU SAEs that modify the activation function to better balance feature detection and strength estimation Rajamanoharan et al. (2024b, a). Some pre-trained repositories, such as Gemma Scope Lieberum et al. (2024) and Llama Scope He et al. (2024), enable broader research.
@@
SAEs Application. SAEs have been used to interpret model representations and understand model capabilities Wu et al. (2025b, c). Beyond static analysis, researchers have begun leveraging SAE-discovered features to steer model behavior. Such activation steering via SAE features has been used to alter attributes like sentiment, truthfulness, or style without fine-tuning the entire model He et al. (2025); Shi et al. (2025); Wang et al. (2025b, a). SAEs have also been applied in the context of model safety and alignment. One study showed that features learned by an SAE from a language model can serve as effective probes for classifying toxic content across languages Bereska and Gavves . By identifying which sparse features correspond to a model’s refusals or safety responses, one can understand and even adjust the model’s safety mechanisms. Intervening on these features has been shown to influence the model’s tendency to refuse or comply with certain prompts Arditi et al. (2024); Yeo et al. (2025); Wu et al. (2025a). Overall, SAEs offer a transparent, feature-level handle on model behaviors that is valuable for safety research.
@@
SAEs Feature Explanation. Inspired by the automated interpretability pipeline that uses GPT-4 to explain GPT-2 neurons from their activating examples (MaxAct) Bills et al. (2023), a framework that has since become the standard for large-scale interpretation of neurons and SAE-learned features in both language and vision models Lin (2023); Huben et al. (2024); Gao et al. (2025). Neuronpedia combines an activation-based method Kissane et al. that highlights the tokens most strongly triggering a feature with a logit-projection method Kissane et al. that infers the feature’s semantic direction by measuring its positive and negative influence on output logits. Recent work proposes an "output-centric" automated feature interpretation that interprets model features not only by considering which inputs activate them, but also by examining the impact of their activation on the model output to generate more accurate and causal interpretations Gur-Arieh et al. (2025).
@@
Agents for Explainability. Recent work has explored using agentic frameworks for explainability. For instance, MAIA Shaham et al. (2024) employs a vision-language model equipped with a set of tools to automate the interpretation of computer vision models. MAIA iteratively designs experiments, composes tools for tasks like input synthesis and exemplar generation, and formulates explanations to explain model behaviors, such as identifying feature selectivity or failure modes. Similarly, KnowThyself Prasai et al. (2026) provides an agentic assistant specifically for LLM interpretability. It unifies various interpretability tools into a single chat-based interface, allowing users to ask natural language questions. In contrast to these applications, our work proposes an agent framework specifically designed to interpret the features learned by SAEs.